OpenStudy (anonymous):

Intro to Trig question The legs of a right triangle measure 5 inches and 7 inches. If theta is the angle between the 5 inch leg and the hypotenuse cos theta ____ A) .81 B) .58 C) .07 D) .71

OpenStudy (aaronq):

 \(sin\theta=\dfrac{5}{x}\) \(cos\theta=\dfrac{7}{x}\) \(tan\theta=\dfrac{5}{7}\) \(x^2=5^2+7^2\) use one of these to solve for the angle